# NeurIT: Pushing the Limit of Neural Inertial Tracking for Indoor Robotic IoT

### v1

@@ VIII Conclusions
In this paper, we present NeurIT, a neural inertial tracking system that takes full advantage of IMU readings. We propose a novel Time-Frequency Block-recurrent Transformer, named TF-BRT, which borrows the idea of time-frequency learning and embraces the advantages of both RNN and Transformer to combine both historical and current motion information. In addition, we leverage the differentiation of the body-frame magnetometers to mitigate the noise introduced by the gyroscope. We implement NeurIT with a customized robotic platform and evaluate it scrupulously across distinct indoor environments. The results show that NeurIT outperforms six state-of-the-art baseline approaches on all benchmarks, achieving a significant accuracy gain of 48% on unseen data compared to the best baseline so far. NeurIT demonstrates remarkable tracking accuracy over extended durations and distances, with a tracking error of one meter over 300 meters, ensuring reliable performance. With the encouraging results, NeurIT takes an important step towards a practical and truly infrastructure-free solution for indoor robotic tracking.

### v2

@@ VIII Conclusions
We introduce NeurIT, a neural inertial tracking system that fully exploits IMU data for reliable indoor localization. At its core is the Time-Frequency Block-recurrent Transformer (TF-BRT), which combines RNN and Transformer with time-frequency learning to capture both long-term dependencies and fine-grained motion patterns. To enhance robustness, NeurIT uses differentiated body-frame magnetometer to compensate for gyroscope drift. We deploy NeurIT on a custom robotic platform and evaluate it extensively. Results show that NeurIT consistently outperforms seven state-of-the-art baselines, achieving up to 48% higher accuracy on unseen data and maintaining an average error of one meter over 300-meter trajectories. These results highlight its strong generalization and real-time capability, marking a significant step toward practical, infrastructure-free indoor robotic tracking.
Acknowledgment
This work is supported by NSFC Grant No. 62222216 and Hong Kong RGC ECS Grant No. 27204522.